What is the probability of randomly picking a P500 bill


1
Expert's answer
2021-12-21T18:32:20-0500

Assume there are 8 P100 bills, 3 P200 bills and 4 P500 bills.  What is the probability of randomly picking a P500 bill

Solution:

No. of P100 bills = 8

No. of P200 bills = 3

No. of P500 bills = 4

Total no. of bills = 8+3+4=15

Now, the probability of randomly picking a P500 bill "=\dfrac{^4C_1}{^{15}C_1}=\dfrac{4}{15}"